OpenStudy (anonymous):

will fan and metal In a city, 49% of the adults are male. 24% of the adults are male and rent action-movie DVDs, and 10% of the adults are female and rent action-movie DVDs. If an adult is randomly selected for a survey, what is the probability that the adult rents action-movie DVDs, given that the adult is female? 0.20 0.37 0.43 0.59 0.15

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

I can help!!......what would do is..do u have a near by calculator

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

ok do 24+10

OpenStudy (anonymous):

34

OpenStudy (anonymous):

so hows that help

Vocaloid (vocaloid):

"the probability that the adult rents action-movie DVDs, given that the adult is female" can be written as P(rents action movie DVDS | female) = P(female and rents action movie DVDs)/P(female) P(female) = 0.51, since 51% of the adults are female (1-0.49 = 0.51) P(female and rents action movie DVDs) = 0.1 (given in the problem)

OpenStudy (anonymous):

ok there is 49% are males...and there are 24% that are renting DVD's ok ad there are 10% of female that are reting movie,,,and they are asking..what is the probability the adult bing a female that is a female you would hik would u try to equal as much as the males o less

OpenStudy (anonymous):

so wat would be the answer?

Vocaloid (vocaloid):

P(rents action movie DVDS | female) = P(female and rents action movie DVDs)/P(female) = 0.1/0.51

OpenStudy (anonymous):

thats not a valid answer they give me

Vocaloid (vocaloid):

divide 0.1 by 0.51 on your calculator...

OpenStudy (anonymous):

no u would do 0.51 /0.1=5.1thn that's the percage

OpenStudy (anonymous):

oh ok

Vocaloid (vocaloid):

0.1/0.51 is about 0.20 when you plug it into your calculator

OpenStudy (anonymous):

i got 5.1

Vocaloid (vocaloid):

other way around... 0.1/0.51

OpenStudy (anonymous):

1.5?

OpenStudy (anonymous):

ohh

Vocaloid (vocaloid):

plug in 0.1, then press the division sign, then press 0.51

OpenStudy (anonymous):

i got 0.1960784314

Vocaloid (vocaloid):

right, so now we round to the nearest hundreths place, and we get 0.20

OpenStudy (anonymous):

so what 1.5 as a percent to u

OpenStudy (anonymous):

thanx @vocaloid

Vocaloid (vocaloid):

no problem, glad I could help :)

OpenStudy (anonymous):

dang it my bad 4 da miss medal

OpenStudy (anonymous):

so i didn't help at least i tried!

OpenStudy (anonymous):

it always seems to be higher level always getting credit ..